OpenStudy (anonymous):

Find F -1(y) F(x)= x-10

OpenStudy (***[isuru]***):

heres how.. lets take y = x - 10 now to find the inverse of y... lets make the "x" subject x = y + 10 now lets switch x and y y = x + 10 heres the inverse function for ya.. f(x) = y = ( x - 10) f^-1(x) = ( x - 10) and the above methods can be used to find the inverse of any function :)